Grade 11Mechanics

A gun kept on a straight horizontal road is used to hit a car traveling along the same road away fron the gun with a uniform speed of 72*21/2 km/h. The car is at a distance of 50m from the gun, when the gun is fired at an angle of 45o with the horizontal. Finda) the distance from the gun from the car when the shell hits it.b)the speed of projection of shell from the gun .[g=10m/s2]

Pihu
9 Years agoGrade 11

1 Answer

somi teez
9 Years ago
hi,     
 u=72*sqrt(2) km/hr=20*sqrt(2) m/sec
now
,      let us consider in time t secods,distance covered by the car=20*sqrt(2)*t
      distance moved by bullet from gun fired=u cos 45*t
we can say,that when the bullet will hit the car distance moved,
= 50+20 sqrt(2) *t=u*(1/sqrt(2))*t..........(1)
since,gum is fired at an angle 45,  so  t=2usin45/g.......(2)
t=2ug/(sqrt 2)
substituting value of t from 2 to 1,  and g=10 m/s
we get the equation   u^2-40u-5=0    then solving quadratic to get u